OpenStudy (anonymous):

Factor the polynomial completely. \[a^4 - b^4\]

OpenStudy (misty1212):

hi!!

OpenStudy (misty1212):

do you know how to factor the difference of two squares?

OpenStudy (anonymous):

Hello! :)

OpenStudy (anonymous):

I think....

OpenStudy (misty1212):

try, i will tell you if you are right

OpenStudy (anonymous):

Well, I am not sure what the first step is... >_<

OpenStudy (anonymous):

Especially with a problem like this.... :(

OpenStudy (xapproachesinfinity):

hint: \[a^4-b^4=(a^2)^2-(b^2)^2\]

OpenStudy (xapproachesinfinity):

now see if that reminds you of something you know

OpenStudy (xapproachesinfinity):

okay let's you do this :)

OpenStudy (anonymous):

So it would be: \((a^2)^2 - (b^2)^2\)

OpenStudy (anonymous):

What is the next step? :)

OpenStudy (xapproachesinfinity):

what is \[a^2-b^2=?\]

OpenStudy (xapproachesinfinity):

Do you recall the difference of two squares?

OpenStudy (anonymous):

Ummmm....

OpenStudy (anonymous):

No...?

OpenStudy (xapproachesinfinity):

hmm eh, how are you supposed to factor without that then lol

OpenStudy (xapproachesinfinity):

okay so we have \[(a-b)(a+b)=a^2-b^2\]

OpenStudy (xapproachesinfinity):

if you expand you will get the right hand side

OpenStudy (xapproachesinfinity):

so \[(a^2)-(b^2)^2=(a^2-b^2)(a^2+b^2)\]

OpenStudy (anonymous):

Oh! I was supposed to say that (a - b)(a + b) when you asked earlier...

OpenStudy (xapproachesinfinity):

of course not, that's just a refresher to you since you don't remember that

OpenStudy (anonymous):

Oh... Okay.

OpenStudy (xapproachesinfinity):

yes of course, blah, you made me worthless talking for minutes for nothing

OpenStudy (xapproachesinfinity):

ok then take what i did and go further then

OpenStudy (anonymous):

So, you are saying that \((a^2)^2 - (b^2)^2\) can be simplified further?

OpenStudy (anonymous):

To something like \((a^2-b^2)(a^2+b^2)?\)

OpenStudy (anonymous):

@xapproachesinfinity ^

OpenStudy (xapproachesinfinity):

(a-b)(a+b)(a^2+b^2) should be this

OpenStudy (xapproachesinfinity):

gotta leave

OpenStudy (anonymous):

Wait, that's the final answer? \[(a-b)(a+b)(a^2+b^2)\]

OpenStudy (anonymous):

\[\Huge ???\]

OpenStudy (anonymous):

\[\Huge \ddot \frown\]